OpenStudy (anonymous):

WILL FAN AND MEDAL!!!!! Solve 2/3n=8

OpenStudy (anonymous):

5.3333333333

OpenStudy (texaschic101):

Is it (2/3)n = 8 or 2/(3n) = 8 ??

OpenStudy (texaschic101):

if it is (2/3)n = 8....then multiply both sides by 3 (3)(2/3)n = (8)(3) 2n = 24 n = 24/2 n = 12 check.. 2/3(12) = 8 24/3 = 8 8 = 8 (correct) n is gonna be 12

OpenStudy (anonymous):

thank you so much and you explained it thats even better

OpenStudy (texaschic101):

YW :)